OpenStudy (anonymous):

Materials expand when heated. Consider a metal rod of length L0 at temperature T0. If the temperature is changed by an amount ∆T, then the rod’s length changes by ∆L=αL0∆T, where α is the thermal expansion coefficient. For steel, α=1.24x10^-5°C. Express length L as a function of T if L0=65 in. at T0=100°C.

OpenStudy (anonymous):

well, first recognize the total length is given by \(L(T)=L_0+\Delta L\)

OpenStudy (anonymous):

we're told \(\Delta L=\alpha L_0\Delta T=\alpha L_0(T-T_0)\) since \(\Delta T=T-T_0\) hence:$$L(T)=L_0+\alpha L_0(T-T_0)$$

OpenStudy (anonymous):

How did you realize that L(T)=L0+ΔL?

OpenStudy (anonymous):

@OddlyWeird the length at temperature \(T\) is the original length \(L_0\) plus the change in length \(\Delta L\) at that temperature